OpenStudy (anonymous):

How would you explain to your friend why the quotient of x^8 and x^2 is x^6 and not x^4 ?

OpenStudy (anonymous):

Rewrite x^8 as x*x*x*x*x*x*x*x/x*x. Simplify using division to get x*x*x*x*x*x which is the same as x^6

OpenStudy (anonymous):

Whi isn't it the same as x^4?

OpenStudy (anonymous):

Because you can't treat exponents like multiplication. While 8x/2x = 4x, x^6/x^2 = x^6. Basically, (x^a)/(x^b) = x^(a-b)

OpenStudy (anonymous):

Is there a much more simple way of writing the x^8 and x^2

OpenStudy (anonymous):

\[x^8 = x*x*x*x*x*x*x*x\] or \[x^2^3\]

OpenStudy (amistre64):

Like terms cancel in fractions: x^8 means you have 8 little "x"s on top part x^2 means you have 2 little "x"s on the bottom part cancel 2 little "x"s from the top and your left with 6 little "x"s. cancel 2 little "x"s from the bottom and your left with the number "1". 6 little "x"s over the number "1" means x^6 :)